OpenStudy (anonymous):

add: 6 9 -- -- + ----- (X+4) (X-3)

OpenStudy (anonymous):

(15x+44)/(x^2-x+12)

OpenStudy (anonymous):

\[\frac{6}{x+4}+\frac{9}{x-3}=\frac{6(x-3)+9(x+4)}{(x+4)(x+3)}\]

OpenStudy (anonymous):

multiply out in the numerator and combine like terms you will get what dheeraj wrote

OpenStudy (anonymous):

or maybe not lets check \[6(x-3)+9(x+4)=6x-18+9x+36\] \[15x-18\]hmm i get \[\frac{15x-18}{(x-3)(x+4)}\]